# Universal thermodynamic topological classes of black holes in perfect fluid dark matter background

### v1

@@ VII Conclusions
The Schwarzschild black hole in PFDM has one state with a negative winding number and, hence, a negative heat capacity. Thus, it belongs to the universal thermodynamical class $W^{1-}$, with thermodynamically unstable small and large black holes at both low and high temperature limits. For the Reissner-Nordström, Kerr, and Kerr-Newman black holes in PFDM background, when $\beta<\beta_{c}$ there is no black hole state whereas for $\beta_{c}<\beta$, there are two black hole states: a small black hole and a large black hole. In the low temperature limit, the small black holes have a winding number of $1$, and the large black holes have a winding number of $-1$, representing thermodynamically stable and unstable states, respectively. Furthermore, at high temperatures, no black hole state exists. The Reissner-Nordström, Kerr, and Kerr-Newman black holes in PFDM background all belong to the same thermodynamical class $W^{0+}$.
Black holes in the AdS background exhibit different topological numbers compared to their counterparts that are not in the AdS background. The Schwarschild-AdS black hole in PFDM background, for low-temperature limit there is no black hole state whereas for high-temperature limit there are two black hole states with stable small size and unstable large size black holes. The Schwarzschild-AdS black hole in PFDM background belong to the universal thermodynamical class $W^{0-}$.
For the Kerr-AdS black hole in the PFDM background, for any chosen values of the black hole parameters, when $\beta<\beta_{a}$ or $\beta_{b}<\beta$, there exists only one black hole state, representing either a stable large or a stable small black hole. For $\beta_{a}<\beta<\beta_{b}$, the Kerr-AdS black hole in the PFDM background exhibits both generation as well as annihilation points, resulting in three black hole states: small, intermediate, and large. At low temperatures, there is a stable small-sized black hole and an unstable intermediate-sized black hole, whereas in the high-temperature limit, there is a stable large-sized black hole and an unstable small-sized black hole. Furthermore, the Kerr-AdS black hole in the PFDM background belongs to the $W^{1+}$ universal thermodynamic class.

### v2

@@ VII Conclusions
The Schwarzschild black hole in the PFDM has one state with a negative winding number and, hence, a negative heat capacity. Thus, it belongs to the universal thermodynamical class $W^{1-}$, with thermodynamically unstable small and large black holes at both low and high temperature limits. For the Reissner-Nordström, Kerr, and Kerr-Newman black holes in PFDM background, when $\beta<\beta_{c}$ there is no black hole state whereas for $\beta_{c}<\beta$, there are two black hole states: a small black hole and a large black hole. In the low temperature limit, the small black holes have a winding number of $1$, and the large black holes have a winding number of $-1$, representing thermodynamically stable and unstable states, respectively. Furthermore, at high temperatures, no black hole state exists. The Reissner-Nordström, Kerr, and Kerr-Newman black holes in PFDM background all belong to the same thermodynamical class $W^{0+}$.
Black holes in the AdS background exhibit different topological numbers compared to their counterparts that are not in the AdS background. The Schwarzschild-AdS black hole in PFDM background, for low-temperature limit, there is no black hole state whereas for high-temperature limit, there are two black hole states with stable-small size and unstable-large size black holes. The Schwarzschild-AdS black hole in PFDM background belong to the universal thermodynamical class $W^{0-}$.
For the Kerr-AdS black hole in the PFDM background, for any chosen values of the black hole parameters, when $\beta<\beta_{a}$ or $\beta_{b}<\beta$, there exists only one black hole state, representing either a stable large or a stable small black hole. For $\beta_{a}<\beta<\beta_{b}$, the Kerr-AdS black hole in the PFDM background exhibits both generation as well as annihilation points, resulting in three black hole states: small, intermediate, and large. At low temperatures, there is a stable-small sized black hole and an unstable-intermediate-sized black hole, whereas in the high-temperature limit, there is a stable-large sized black hole and an unstable-small sized black hole. Furthermore, the Kerr-AdS black hole in the PFDM background belongs to the $W^{1+}$ universal thermodynamic class.
This study can be summarized as follows: the universal topological classes and the stability analysis based on these classes of black holes with PFDM background are consistent with those of black holes without PFDM background, indicating that PFDM does not alter the results rtop3unipaper ; rtop4unitopo .